OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. x2 - 16 = 0 and x2 = -2x + 24 Part 2: Describe what the solution(s) represent to the graph of each. Part 3: How are the graphs alike? How are they different?

OpenStudy (anonymous):

MEDALS FOR EVERYONE!!!

OpenStudy (campbell_st):

1st quadratic is the difference f 2 squares (x-4)(x+4) then x = -4, 4 this is concave up... symmetric about the y axis and has y intecept y = -16 and x intercepts at -4 and 4  2nd is x^2 +2x - 24 = 0 factors of -24 that add to 2 (6 and -4) so it factorises to x^2 + 2x -24 = (x + 6)(x-4) x = -6. 4 graph is concave up, symmetric about the line x = -1 has a y intercept at y = -24 x intercepts at x = -6 and 4 has a minumum value at (-1, -25)